OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (anonymous):

If one line is parallel to another, then they have the same slope. Therefore you need to figure out the slope of the line. Since you have two points, you can calculate this. Do you know how?

OpenStudy (anonymous):

no, i came into this class halfway through the semester and missed it

OpenStudy (jackellyn):

y=mx+b where m is the slope. m= y2-y1/x2-x1 y2=-4 y1=6 x2=-1 x1=4 Can you do it now?

OpenStudy (jackellyn):

@StevenReynolds